OpenStudy (anonymous):

How to solve 2y+4x+8;x=-2,1,3 PLease tell me step by step!!!

OpenStudy (chihiroasleaf):

are you sure that it's the question? it's not an equation...

OpenStudy (anonymous):

Like if you take off the ;x=-2,1,3 it has the 2y+4x+8;x

OpenStudy (anonymous):

that is how they have it in the book so i dont know.

hartnn (hartnn):

2y+4x+8 is an expression. you cannot solve an expression ,only find its value at a particular 'x' and 'y'. did you mean 2y = 4x+8 ?? and need to find the value of y for x= -2,1,3 ?? if possible , you can post the screenshot of the entire question, so we can help :)

OpenStudy (anonymous):

Ok ill try

OpenStudy (anonymous):

Just to let yall know that equation falls under Rewriting a literal Equation it wont let me put pic sorri

OpenStudy (anonymous):

The way it is in the book is like this 2y+4x=8;x=-2,1,3

hartnn (hartnn):

ok,so now i think you need to find 'y' values. can you isolate 'y' from 2y+4x= 8 ? that is bring it in the form y=....

OpenStudy (anonymous):

Nope im still confused

hartnn (hartnn):

2y+4x=8 divide entire equation by 2, what u get ?

OpenStudy (anonymous):

one sec let me see

OpenStudy (anonymous):

i got 4

OpenStudy (anonymous):

So isnt it going to be 4x=8-2y?

hartnn (hartnn):

let me give u an example, if you had 6y+2x=10 and if you wanted to divide entire equation by 2, you would get , (6/2)y + (2/2)x =10/2 which is 3y+x=5 , an equation. so what do u get for 2y+4x=8 ?

OpenStudy (anonymous):

So isnt it going to be 4x=8-2y? then 4x/4=(8-2y)/4 then answer is x=(8-2y)/4

hartnn (hartnn):

but here, you need to isolate y, you actually isolated 'x'. can you try to isolate 'y' similarly ?

OpenStudy (anonymous):

Nope i just know how to do the 'x'

hartnn (hartnn):

thats why i asked you to divide both sides by 2 first and gave you an example, see that and tell me what u get after dividing 2 on both sides of 2y+4x = 8

hartnn (hartnn):

alternatively, you can subtract 4x from both sides first...

OpenStudy (koikkara):

2y + 4(-2) = 8 2y + (-8) = 8 2y = 8+8 y = 16/2 y = 8 ---------------------------- 2y + 4(1) = 8 2y + 4 = 8 2y = 8-4 y = 4/2 y = 2 ------------------------------ 2y + 4(3) = 8 2y + 12 = 8 2y = 12-8 y = (-4)\2 y = -2

OpenStudy (koikkara):

sorry to say...just substitute it....for x....

OpenStudy (anonymous):

Ok thank you

OpenStudy (koikkara):

wlcm..